OpenStudy (anonymous):

d/dx (ln(x+1))/(ln(x-3))

OpenStudy (anonymous):

Hint: Dividing log functions is the same thing as subtracting them, that should make things easier.

OpenStudy (anonymous):

honestly i just want to know if the quotient rule will get me ln(x-3)^3 or (ln(x-3))^2

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

I don't feel like typing it out so i'll let someone else do it.

OpenStudy (anonymous):

sorry i didnt word that right im talking about the bottom of the fraction made by the quotient rule

OpenStudy (anonymous):

everything will be over the bottom term squared.

OpenStudy (anonymous):

Simplifying you get: ln((x+1))-(x-3)) can you solve that?

OpenStudy (anonymous):

yes but would the bottom term squared be \[\ln (x-3)^{2} or \ln ^{2}(x-3)\]

OpenStudy (anonymous):

You should get 1/(x+1) - 1/(x-3)

OpenStudy (anonymous):

brainshot im taking the dreivative of a function not simply dividing those numbers thats what the d/dx is for

OpenStudy (anonymous):

You are forgetting the rules for taking a derivative of a ln function. it is 1/(original equation) * derivative of the original equation. In this case ln(x-3) = Derivative of (x-3)/The original equation, so you will get: 1/(x-3)

OpenStudy (anonymous):

okay sorry i didnt even know that rule i was trying to use the quotient then chain rule thank you!